# Title keyword 5 for citation Title for citation Title for citation

### v2

@@ 3 AI for Network
3.1 Definitions of AI4Network
AI for Network is the notation of using AI to improve the performance, efficiency and user service experience of the network itself. The primary research of AI for NET includes using AI to optimize traditional algorithms, optimize network functions, optimize network operation and maintenance management, etc., to improve the transmission efficiency and wireless performance of the wireless air interface and network sides. Figure 7 illustrates a schematic of AI’s capabilities for the network.
AI enables wireless networks through operation and maintenance efficiency improvement, network performance enhancement, and new business enabled. AI can allow wireless networks based on AI’s capabilities, including feature extraction, Forecast, Self-adaptation, Optimization, Real-time, correlation, Scene clustering and other capabilities. These capabilities are the basis for AI to enable wireless networks. In the user’s environment, AI optimizes traditional methods to improve network performance on various terminal devices. AI leverages its ability to assist in optimizing the network, and different cases reflect AI’s capabilities. AI can optimize network element functions and central management in the upper-layer network to improve network operation and maintenance efficiency. At the same time, the generation of AI has also spawned a series of new services to improve and better serve users.
@@ 3.1 Definitions of AI4Network
3.2 Improved O&M Efficiency
AI can analyze large amounts of data and make decisions in real time [73]. This can be particularly useful in wireless networks, where many variables and parameters must be constantly monitored and adjusted to optimize the system’s performance. Table 5 lists the wireless use cases of AI.
AI algorithms can analyze data from the wireless network to predict when maintenance is needed, allowing maintenance to be performed proactively and timely. It can dynamically allocate resources, such as bandwidth and power, in the wireless network to optimize the system’s performance.
The air interface data is affected by the environment and has significant time variability, and the data quality is easily affected by many factors. Such sparse, high-dimensional data is very suitable for building a mapping relationship between the input and output of AI models, and the wireless data generated at all times ensures the number of samples for model training.
Comparing AI technology with traditional air interface technology, we show the unique advantages of AI itself and point out the development direction of the deep integration of wireless air interface technology and AI in the future. It also signifies the beginning of the communication network’s evolution toward increased intelligence and autonomy.
Table 5: Wireless use cases of AI.
Classification Use Cases & AI capability
Improved O&M Efficiency Traffic Prediction Algorithm & Forecast
Network Parameter Optimization & Optimization
Digital Twin
Air Interface Performance Enhancement CSI Feedback & Feature extraction
OFDM Receiver & Self-adaptation
Beam Management Algorithm & Feature extraction
BS Energy Conservation & Forecast
New Business Enablement Mobility Management & Forecast
Network Slicing & Optimization
Wireless Localization Algorithm & Real-time
Environment Awareness & Real-time
@@ 3.2.1 AI-Based Traffic Prediction
With the wide application of 5G communication, users’ demand for data traffic continues to increase, and traffic prediction is also paid more attention in communication, as shown in Figure 8. Considering the increasing network load, analysis data based on network traffic can effectively reconfigure the structure of multi-provider networks. Traffic prediction is a means[74] for various networks to realize intelligent network operations such as resource management and predictive control. A forecasting method based on historical averages can be used in traditional traffic forecasting. This approach is based on the assumption that network traffic has a cyclical nature. By averaging the historical data, it is possible to predict the traffic at the next moment.
Figure 8: The structure of traffic prediction
However, traffic forecasting methods based on historical averages have specific bottlenecks. Although the model of this method is simple, it lacks the use of geographic location, data change trend and other information, resulting in obvious errors. The traditional traffic prediction method makes fitting the complex time-varying communication traffic model challenging. In this case, introducing AI technology has become a new direction to solve this problem. At present, AI algorithms have been widely used in the field of mobile communication traffic prediction. For example, algorithm models based on machine learning include Random Forest (RF), Multiple Linear Regression, Gaussian Process Regression, and Prophet time series prediction algorithm. These machine learning algorithms perform better than traditional statistical models but still struggle to capture the spatio-temporal characteristics of the data. Deep learning can effectively extract intrinsic features, such as continuity and timing of traffic information. Compared with traditional statistical models, deep learning models have strong adaptability when dealing with massive and complex communication traffic data. They can extract multi-level features and adjust themselves in time.
Examples such as Residual Networks (ResNet), 3D Convolutional Networks (3D-CNN), GAN and other deep learning models can be used. These models can analyze user behavior patterns and device characteristics, learn the internal relationship between user data to accurately predict user demand for different services and applications, and optimize the resource allocation of the network based on this. At the same time, these models can also analyze network traffic data, identify factors that may lead to network failures, and predict the occurrence and recovery of failures, helping operators diagnose and solve problems faster and accurately and improve network reliability and stability.
In [75], random forest and linear regression are combined to solve the traffic prediction problem in backbone optical networks. In [76], a network traffic prediction model based on synchronous Periodic LSTM (DSP-LSTM) was proposed, which can comprehensively consider the good fitting ability of a linear model to periodic characteristic data and the nonlinear prediction characteristics of neural networks. [77] The internal mechanism of network traffic outburst is explained by extracting the time series of the newly generated number of network flows. The number of network flow-based traffic prediction models shows noticeable performance improvement over the original LSTM and TCN models. In [78], a hybrid prediction model for passive optical network traffic was proposed, which combines fully integrated empirical mode decomposition with adaptive noise with particle swarm optimization support vector regression and LSTM neural networks, showing high prediction accuracy. In [79], a network traffic prediction method based on time representation and spatial convolution was proposed. Graph convolution network was used to explore the topological features of network nodes, and LSTM was used to characterize the time features of the network to help the network learn daily and weekly long-term time features.
@@ 3.2.2 AI-Based BS Energy Conservation
The traditional network energy-saving algorithm usually considers the current load of the BS to control the opening and closing of the BS [80, 81]. Traditional energy-saving strategies mainly involve device-level and site-level energy conservation [82]. At the device level, hardware energy-saving schemes are primarily studied from device and hardware design, and some RF power amplifiers are turned off when the BS load is low. At the site level, software energy-saving schemes are mainly studied from a symbol (subframe), carrier, channel turn-off, deep hibernation, etc. By changing the BS mode at a low load, the effect of energy conservation is achieved.
Nowadays, mobile communication networks are large and complex, requiring high-precision real-time data processing and system-level coordination and optimization capabilities. The traditional BS energy-saving algorithm has obvious defects in data processing, operations optimization and system comprehensive optimization. Therefore, the current energy conservation of BSs includes not only equipment level and site level energy conservation but also network level energy conservation. Figure 9 illustrates the structure of BS energy conservation. The blue line indicates the traffic required by the base station, and the thick black line indicates the actual traffic consumption. It can be seen that the actual traffic consumption far exceeds the exact needs of the base station, and the thin black line makes the traffic consumption we want to see. From the perspective of multi-network coordination, intelligent energy conservation scheme research is carried out, and AI strategy deployment plays a crucial role in network-level energy conservation [83].
Figure 9: The blue line indicates the actual flow of the BS, and the bold black line indicates the actual flow of the base station, which shows that the actual flow is much higher than the needs of the BS, so we hope to meet the traffic demand and reduce the energy consumption of the BS like the thin black line.
AI strategy can quickly analyze large-scale energy consumption data and discover hidden rules and correlations [84]. Therefore, network traffic can be accurately predicted according to dynamic data analysis results to reasonably plan parameters such as BS power and hibernation state and achieve accurate energy consumption control and resource optimization. The AI method can also systematically consider many factors, such as network coverage, capacity, and signal quality, to dynamically adjust the energy-saving strategy in the network and realize the global optimization decision. In addition, primary data such as the configuration and performance statistics of the live network can also be used to optimize the 5G device-level energy-saving function parameters based on feature clustering and neural network algorithms. Under the premise of ensuring service quality, the problem of 4G/5G overlapping coverage can be fully considered for the cell’s cooperative shutdown to maximize the effect of the device-level energy conservation function [85].

### v3

@@ 2 AI for Network
The most fundamental goal of the integration of wireless communication and AI lies in enhancing the performance, efficiency of the network itself and the service experience of users by virtue of AI, which is known as AI4NET [36]. At this stage, the focus is on how to utilize AI algorithms to optimize communication algorithms and network functions. For example, AI can be used to optimize the signal modulation and demodulation process, making signal transmission more accurate and efficient; with the help of AI, network resources can be intelligently allocated to achieve load balancing; and an automated O&M management mode can be constructed to improve O&M efficiency and reduce costs. Under this mode, the introduction of AI will not have a significant impact on the original network architecture. It mainly applies and improves specific identified communication problems by training AI algorithm models, just like precisely implanting intelligent patches into the network to gradually improve the intelligence level of the network while maintaining the stability of the architecture.
The conceptual diagram of AI4NET is presented in Figure 5, where the wireless AI possesses capabilities such as feature extraction, prediction, adaptation, optimization, real-time processing, correlation, and scene clustering. These capabilities form the foundation for AI-enabled wireless networks and can be directly deployed on the BS side and the CN to empower wireless communication. In the user environment, AI optimizes conventional methods to enhance network performance on various terminal devices. On the network side, AI takes advantage of its capabilities to improve the quality of end-to-end services, optimize the functions of network elements, conduct automated management of the upper-layer network, and boost network O&M efficiency. Meanwhile, the advent of AI has also given rise to a series of new services aimed at improving and serving users in a better manner.
AI4NET has initiated relevant research and applications in 5G. In 6G, with the more mature AI technologies represented by DL and the emergence of new infrastructures integrating connectivity and computing power, the relevant applications will become more abundant and sophisticated and may further evolve in depth. Next, this section will enumerate some cases from two aspects, namely air interface performance enhancement and network O&M improvement, to better illustrate the concept of AI4NET.
@@
2.1 Air Interface Performance Enhancement
The performance of the wireless air interface will directly affect the user’s communication experience. Finding a better air interface algorithm to optimize air interface performance has always been the goal of industry insiders. However, user demands are increasing with the continuous increase of communication services. conventional algorithms incur complexity through complex matrix operations and reach the performance bottleneck.
The booming development of AI has provided a new direction for air interface transmission in communication. AI can fit complex and nonlinear mapping relationships. Compared with conventional methods, AI is more suitable for applications in complex time-varying wireless environments. Thus, AI can replace conventional communication methods to improve wireless network performance significantly. In this section, we show how AI technology can improve the transmission performance in CSI feedback, orthogonal frequency division multiplexing (OFDM) receivers, beam management, and wireless localization.
2.1.1 AI-Based CSI Feedback Algorithm
Within the framework of 6G communication technology, the feedback overhead experiences a sharp increase along with the significant rise in the number of antennas. Due to the extremely high demand of 6G networks for refined resource utilization, the uplink resources become increasingly scarce and precious. Consequently, ensuring the accuracy of channel reconstruction and minimizing the CSI feedback cost is one of the bottleneck problems that need to be surmounted. The utilization of AI technology in CSI feedback is capable of profoundly dissecting these elaborate channel characteristics. Through training with a substantial quantity of labeled samples, the AI model can acquire the characteristic patterns of the channel within diverse fading environments, interference scenarios, and multipath circumstances. In contrast to conventional CSI estimation approaches relying on statistical theory or simple linear models, AI-based methods exhibit stronger generalization competencies and can estimate CSI with greater precision in complex real-world communication settings, effectively diminishing the estimation error. This highly accurate CSI estimation facilitates the communication system to fine-tune transmission parameters, such as modulation and coding schemes and transmission power, more accurately, thus augmenting the spectral efficiency and transmission reliability of the system.
The massive CSI data and the inherent random characteristics of CSI make it preferable to introduce AI for designing a new CSI feedback mechanism. For instance, the nonlinear characteristics of DL can be utilized to efficiently extract the features of CSI, and the original CSI data can be transformed into a more compact feature representation, thereby reducing the communication overhead in the CSI feedback process. An example of its implementation is the autoencoder architecture, which employs neural networks to extract and compress the features of CSI channel information. The authors of [37] were the first to apply DL to CSI feedback first. The proposed CsiNet scheme utilizes convolution to extract channel features, compress and reconstruct channels, and it is superior to the CSI feedback scheme based on compressed sensing in terms of feedback accuracy and computational complexity. The encoder of CsiNet compiles the high-dimensional CSI into code words, while the decoder is responsible for decoding the code words back into the original CSI. The internal structures of encoders and decoders of different algorithms vary, resulting in differences in channel reconstruction performance. As shown in Figure 6, Huawei Technologies Co., Ltd. also demonstrated the huge advantages of AI-based CSI feedback over conventional codebook approach through system-level simulations and AI with additional information can effectively improve the performance on the square of the generalized cosine similarity (SGCS) and throughput [38].
(a) SGCS versus bits
(b) Throughput gain versus bits
Figure 6: The performance of R16 Type II codebook, AI-SF and AI-SFT on SGCS and throughput gain. AI-SF means the AI/ML-based spatial-frequency compression method, which transformer is used as the backbone of both encoder and decoder. AI-SFT is AI-SF with additional past CSI.
The studies in [39] increased the size of the convolutional kernel based on CsiNet to improve the perceptual field of view of the convolutional layer, which was conducive to extracting the sparsity of the channel and further enhancing the accuracy of channel reconstruction. Although a larger perceptive field of view could effectively extract the sparsity of CSI, a smaller convolution kernel could extract finer features from CSI. A DL-based CSI compression and quantization method was developed in [40] considering high compression ratios. A multi-rate CSI compression framework was designed in [41] to improve the generalization of the model in the field of transfer learning. In [42], a non-local neural network was introduced based on the CsiNet network to capture a wide range of dependencies, and the accuracy of channel recovery was improved compared to CsiNet. The CRNet was proposed in [43], which used convolution kernels of different sizes for channel feature extraction and recovery, reducing the computation and improving the accuracy of channel reconstruction.
The above-mentioned literature aimed at using DL to improve the compression and recovery accuracy of CSI. However, in order to better deploy the AI-based CSI feedback model in the actual wireless environment, model monitoring, model updating and AI-based signaling process management are hot issues that require our attention in future studies.
2.1.2 AI-Based OFDM Receiver
The wireless communication transmitter and receiver are divided into multiple processing mods in OFDM. Each module is responsible for specific sub-tasks, such as Quadrature amplitude modulation, an inverse fast fourier transform, channel estimation, signal detection, etc. However, conventional receiver often relies on pre-set mathematical models and assumptions. It works well under ideal assumptions, but in practical applications, the channel conditions are complex and changeable (such as multi-path propagation, frequency shift, nonlinear distortion, etc.), and conventional receivers may not be able to effectively cope with these changes especially for channel estimation module.
AI method can extract essential information in sparse and time-varying pilot frequencies. The complex mapping relationship between the input and the output is effectively constructed, thus AI-based receiver can learn and adapt to channel characteristics in complex environments. Each module of the conventional receiver requires accurate signal modeling and calculation, which leads to the complexity and calculation burden of the system. The AI model is able to extract features directly from the original signal and demodulate it in a data-driven manner, simplifying the complex process of conventional receivers.
Figure 7: This is the difference between conventional OFDM and AI-based OFDM receivers. Using AI modules to replace conventional physical layer processing, AI’s capabilities can compensate for signal distortion.
Regarding an OFDM wireless receiver, some researchers have used a fully connected deep neural network (DNN) to improve the existing modular OFDM receiver. A typical framework of AI-based receiver is given in Figure 7. We use DNN parameters $\Phi$ to construct the relationship between input $X$ and output $Y$ as $Y:f(X,\Phi)$. With the assumption of $L$ hidden layers, we denote $w_{l}$ as the parameter weights of $l-th$ layer, where $l$ ranges from 00 to $L+1$. $w_{0}$, and $w_{L+1}$ represent the weights of input layer, output layer, respectively. $b$ denotes the bias of each layer. In the training stage, the optimization direction of model is minimizing the loss between the output $Y$ and true label $Y^{true}$. Thus the network parameters $\Phi$ are updated by changing the $w_{l}$ during the process of back propagation. Hence, the AI model can encode the information of actual scenes into neural networks and optimize the performance of algorithms by adjusting the structure and parameters of neural networks. Thus, the performance of the receiver is significantly improved.
In [44], a model-driven DL approach was proposed, which combines DL with expert knowledge to replace existing OFDM receivers in wireless communications. An ML-assisted physical layer receiver technique was proposed to demodulate the OFDM signals, which were subject to very high Doppler effects and corresponding distortions in the received signal [45]. To find a balance between full-size cyclic prefix (CP) and non-existent CP, the authors of [46] investigated the redundancy problem and proposed a minimum redundant OFDM receiver using DL tools. In [47], the receivers were designed based on DNN in DL, consisting of a layer of DNNs and soft decisions. The problems of channel estimation error, delay, and decoding limits between users with conventional detection methods have been solved. It can recover the symbols of all users at one time and jointly perform channel estimation and signal detection. A novel generative supervised DNN was designed in [48], which could be trained using a reasonable number of pilots. After channel estimation, the neural network-based receiver (specifically, the autoencoder) jointly learns the pre-encoder and decoder for data symbol detection. An intelligent receiver for the OFDM communication system was designed in [49] based on the DNN structure, and the intelligent receiver was realized by optimizing the DNN structure. This method can recover information at the receiving side and avoid complicated pilot operations and signal error accumulation. Moreover, a convolutional neural networks (CNN) was used to reduce the bit error rate by using the mathematical calculation function of discrete Fourier transform and the training of OFDM signal samples [50].
Further, when we replace the transmitter with AI module and train the transmitter and receiver jointly, we can realize end-to-end communication. However, in existing studies, the channel model in end-to-end communication is simple, and the existing methods are difficult to directly extend to actual deployment. Meanwhile, it is still necessary to improve the model’s generalization ability for complex wireless channels. For practical deployment, we should take into account the difficulty of solving the gradient direction propagation of online training. AI-based receivers have shown great benefits and development potential in various conditions. For actual deployment, we still need to further consider the model’s generalization ability on potential new environments, as well as its robustness in harsh environments.
2.1.3 AI-Based Beam Management
In light of the necessity for 6G networks to support elevated data rates, diminished latencies, and broader connectivity, beam management technology assumes a pivotal function in the efficient exploitation of spectrum resources, the augmentation of communication coverage, and the assurance of communication quality. The massive multi-antenna system within 6G engenders a greater quantity of beams with more constricted widths, consequently encountering novel challenges with respect to beam alignment.
In the existing academic literature, the beam selection methods of multi-antenna systems can be divided into several categories, including beam selection based on beam scanning [51], beam selection based on training signals [52], beam selection based on position prediction [53], and beam selection based on hierarchical search [54]. However, the conventional beam management methods have some limitations when applied to millimeter wave massive multiple-input multiple-output (MIMO) systems [55]. The primary cause is that the real-world wireless propagation environment often has various characteristics. After a beam is narrowed, the signal transmission is more susceptible to occlusion. The increase in the number of beams significantly increases the number of beam search operations in beam alignment. It is difficult to accurately select the optimal beam instantaneously.
With the application of AI, the optimal beam can be selected without scanning all the beam pairs, thus alleviating the scanning overhead existing in the conventional beam management method as shown in Figure 8. In [56], DNN was utilized for beam alignment based on the context information of the UE location. The proposed method can reduce the search time by 4x for AP selection and over 10x for beam selection. By taking advantage of the mmWave channel structure, a novel hierarchical beam alignment framework was proposed in [57] to leverage DL techniques to seek the optimal beamformer with learnable probing codebooks. This novel framework learns two tiers of probing codebooks and uses their measurements to predict the optimal beam in a coarse-to-fine search manner. A DL-based beam alignment method was proposed in [58] by jointly training the probing codebook and the beam predictor. Simulation results demonstrate that the proposed method can achieve high beam alignment accuracy while reducing the beam sweeping complexity by 10×.
The above literature aims to use DL methods to improve the accuracy and robustness of beam alignment, and has achieved excellent performance, demonstrating the superiority of AI. For practical deployment, we should further consider the acquisition of high-quality tags in complex environments. Considering the training overhead, improving the generalization ability of AI models among different BS environments also becomes an important future work.
In the actual communication scenario, the massive historical beams provide enough data for training AI models, and the search for the optimal beam based on the correlation between beams constructed by AI dramatically reduces the search space of the beam. At the same time, due to the implicit structure of AI, more complex functional relationships can be explored. Compared with the conventional beam management method received signal strength (RSS), even in a low signal-to-noise ratio (SNR) environment, the optimal beam selection based on AI still has obvious advantages and robustness.
2.1.4 AI-Based Wireless Localization
The 6G network is mandated to furnish not only ultra-high-speed data transmission and ultra-low-latency communication services but also high-precision wireless location capabilities for diverse application scenarios. Ranging from the accurate positioning of vehicles in intelligent transportation systems, to the location tracking of devices in the industrial IoT, and to indoor high-precision location services, wireless location represents one of the crucial technologies for actualizing these applications. Wireless positioning of 6G refers to obtaining the location information of mobile users in a cellular environment. This location information is usually in the form of the mobile user’s geographic coordinates relative to some reference points. Wireless location is also commonly referred to as mobile location, wireless location, and geolocation [59].
Wireless positioning technology can be divided into wide and local area positioning according to the scope of services. Wide area positioning provides users with universal positioning services. Local area positioning is mainly used in indoor scenarios. Wide area positioning primarily includes mobile communication network positioning and global satellite navigation systems, e.g., the global positioning system (GPS). The GPS is a kind of high-precision radio navigation positioning system based on artificial Earth satellites, which can provide accurate geographical position, vehicle speed and accurate time information anywhere on the Earth and in near-earth space. Local wireless positioning mainly includes Wi-Fi, Bluetooth, UWB, etc. Its positioning technology can be divided into four types: ranging positioning based on RSS, fingerprint positioning based on signal characteristics, positioning based on signal transmission time and angle of arrival measurement.
Many applications need high-precision positioning, such as industrial automated guided vehicle and asset tracking, especially indoor precision positioning, but GPS cannot be used indoors. Despite having a specific positioning reference signal, LTE cannot meet the requirements for high-precision positioning due to its relatively low positioning accuracy and the large distance between BS (about 100 meters even with a 20 MHz bandwidth). Bluetooth, Wi-Fi and other wireless location-based technologies have high deployment costs and are difficult to become a universal positioning technology [60].
The application of AI technology has been introduced to achieve high-precision positioning in the 5G era. One approach is to utilize the random forest technique [61], which can create a classification model to split a vast area into many small grids and then predict the grid where the user is located. It can also make a regression model to predict the user’s position coordinates. In another method, multi-layer perceptron has a similar application to random forest and can establish regression and classification models to locate targets [62]. AI algorithms can also be trained to recognize the fingerprint of a wireless signal in a specific location and process and fuse data from different sensors to achieve high-precision positioning.
[…]
2.2 Network O&M Efficiency Improvement
For the continuous improvement and optimization of the performance of mobile communication networks, the introduction of AI has also played a crucial role in the relevant fields on the wireless communication network side, providing brand-new solutions to the problems in key areas such as data traffic prediction, BS energy-saving management, network parameter optimization, mobility management, and Resource Orchestration. AI can analyze a large amount of data and make real-time decisions, which is especially useful in wireless networks where many variables and parameters must be continuously monitored and adjusted to optimize system performance [64]. It can dynamically allocate resources such as bandwidth and power in wireless networks to optimize system performance.
In addition, the integration of AI enables the network to better understand and predict the complexity and dynamics of the network. Based on the prediction results, proactive and timely network maintenance can be carried out. Furthermore, by leveraging AI’s DL and pattern recognition technologies, a more in-depth understanding of network behavior can be achieved, realizing more optimal resource management and scheduling strategies. Subsequently, the role of AI in network O&M management will be introduced through specific cases.
@@ 2.2.1 AI-Based Traffic Prediction
The progressive deployment of 6G networks [65] heralds the increasing prevalence of new application scenarios like virtual reality (VR) and immersive communication. In such applications, the network traffic exhibits high variability and significant dynamics. In this context, precise and efficient network traffic prediction techniques have become particularly important. By predicting network traffic, it is possible to identify traffic fluctuations and peak periods in advance, thereby enabling dynamic allocation and scheduling of network resources.
In recent years, data-driven DL methods have gradually become the mainstream technology in the field of traffic prediction. These methods, leveraging the powerful modeling capabilities of DNNs, can effectively extract and capture potential feature relationships from complex and dynamically changing network environments. For instance, the authors of [66] employed recurrent neural network (RNN) to achieve joint spatiotemporal prediction, aiming to improve the accurate modeling of network traffic variations. The method proposed in [67], which is based on deep belief networks, effectively predicts the long-term variation trends of network traffic. The result in [68] demonstrated that, compared to conventional ̵autoregressive integrated moving average models, long short-yerm memory outperforms in traffic prediction tasks, particularly in handling nonlinear and complex time-series data.
In centralized traffic prediction methods, BS typically need to collect traffic data from various geographical locations, which inevitably introduces additional communication overhead and potential security risks. As a result, federated learning (FL) based traffic prediction methods have emerged. The FedDA method was proposed in [69] effectively reduces the delay and bandwidth overhead associated with data transmission by only transmitting model parameters instead of raw traffic data. This approach not only ensures privacy and security but also improves the efficiency of data processing. As shown in Figure 9, the authors of [70] proposed a distance-weighted federated traffic prediction method that better captures the spatiotemporal characteristics of traffic and improves the accuracy of traffic forecasting.
To reduce the model training time, the authors of [71] proposed a transfer learning-based approach, which accelerates the training process by uncovering the potential correlations between traffic patterns in different regions. Furthermore, other studies have employed techniques such as data sharing [72] and meta-learning [73] to further optimize the performance of traffic prediction models. In the future, network traffic prediction methods will focus not only on extracting spatiotemporal features and long-term dependencies but also on enhancing the lightweight design and privacy protection capabilities of the models.
Figure 9: Network traffic prediction
@@ 2.2.2 AI-Based BS Energy Conservation
According to statistics, the information and communication technology (ICT) industry directly contributes approximately 4% of global greenhouse gas emissions, with mobile communication networks accounting for over 10% of this figure. In response to the challenges posed by climate change, the ITU released Recommendation 1470, titled “Greenhouse Gas Emission Trajectories for the ICT Industry in Line with the United Nations Framework Convention on Climate Change Paris Agreement,” which calls for mobile operators to achieve a 45% reduction in carbon emissions from 2020 to 2030, aiming to drive the industry’s transition towards sustainability. Currently, the energy consumption of communication networks is primarily concentrated in computers and data centers, BS, and related network infrastructure, making energy-saving a critical issue that requires urgent attention.
The energy optimization within 6G networks exhibits a close correlation with user latency. A key focus of academic research has been how to achieve network energy savings by effectively shutting down BS while ensuring the maintenance of basic service quality. The authors of [74] proposed a BS shutdown strategy based on traffic thresholds. However, during peak network periods, this approach may result in frequent state switching, leading to service interruptions. On the other hand, the authors of [75] classified BS states into four different sleep modes based on the activation time scale of the BS. While deeper sleep modes can significantly reduce energy consumption, they also result in increased wake-up delays, which may affect service quality.
The energy-saving strategy for BS predicated on RL, via the incessant interactive learning between the agent and the operational milieu, dynamically modifies the parameters so as to accommodate the variations in service load and the oscillations in channel conditions. In [76], a BS state management scheme based on Q-learning was proposed. This approach dynamically adjusts the energy consumption of the BS in response to load changes by exploring the optimal duration of different sleep modes. The authors of [77] considered the interference issue between multiple BS and designs an energy optimization method based on collaborative strategies. By enabling cooperation between BS, this method reduces the overall system energy consumption. A cascading RL method was proposed in [78], which optimizes the decision process of user management and BS states in multi-BS scenarios, further improving network efficiency. The aforementioned studies primarily focus on internal strategy optimization within communication networks, with insufficient attention paid to the impact of external resources.
Introducing renewable energy, especially clean energy sources such as solar power, as the energy supply for networks can significantly improve network energy efficiency and reduce dependence on conventional power grids. As depicted in Figure 10, the authors integrate RL with graph neural networks and accomplish the optimization of energy conservation and carbon emissions in wireless networks by capitalizing on photovoltaic solar technology [79]. Looking ahead, the energy-saving efforts in wireless networks should be deeply integrated with the use of renewable energy, providing more sustainable solutions.
Figure 10: Carbon emissions of different strategies